OpenStudy (anonymous):

how do you solve: subtract -5/10-(-3/5)=

OpenStudy (anonymous):

first you need to find the common denominator

OpenStudy (anonymous):

Which is 10

OpenStudy (anonymous):

yes, so what do you have to do to get to the second fraction on the denominator

OpenStudy (anonymous):

-5/10-(-6/10) My answer is 1/10 I'm just not for sure if it should be negative or not

OpenStudy (anonymous):

You have to distribute the negative into the parenthesis

OpenStudy (anonymous):

that is correct and it is should be positive \[\frac{ -5 }{ 10 } + \frac{ 6 }{ 10 } = \frac{ 1 }{ 10 }\]

OpenStudy (anonymous):

the positive is greater than the negative so, the answer would be positive when you add them

OpenStudy (anonymous):

so the two negatives did cancel each other out. Okay thanks I wasn't for sure since it was subtraction.

OpenStudy (anonymous):

not really negative * negative = positive that is where the positive sign came from

OpenStudy (anonymous):

Okay gotcha! Thank you!!

OpenStudy (anonymous):

No problem and welcome to OpenStudy